# Expanding the reach of diffusing wave spectroscopy and tracer bead microrheology

### v1

@@ II.2 Analysis of the short-time correlation function decay without bias
For $t\to 0$ the ICF tends to a constant value of $g_{2}(0)-1=\beta$ where $\beta\lesssim 1$ denotes the coherence factor of the detection fiber optics. In a typical DWS microrheology experiment, for short relaxation times up to some ten or hundreds of microseconds, the bead motion often leads to relatively minor decays of the measured ICF, often only by a few per cent. Extracting $\Delta\vec{r}^{2}(t)\right>$ from these minor decays is strongly affected by the accuracy with which one can determine $\beta$ and by the significant photon shot noise at short correlation lag times [45]. The normalization to determine $\beta$ is usually performed by extrapolating the measured $g_{2}(t)$ to $t=0$ using a linear or polynomial fitting function. Once $\beta$ is known, the normalized ICF is used to determine e.g. $\Delta\vec{r}^{2}(t)\right>$. This standard procedure suffers from two problems. First, the choice of the polynomial order used for the fit, as well as the time interval used for fitting the data, influences the exact value of $\beta$ that is obtained, thereby introducing bias. Here, it is important to note that, due to the inertia of the beads and the fluid, the short-time dynamics are rather complex even for simple colloidal suspensions in water, and thus no simple temporal scaling for the actual $\Delta\vec{r}^{2}(t)\right>$ can be assumed [31, 20, 46]. Second, since the raw ICF is only divided by $\beta$ but otherwise used as-is for further analysis, the experimental noise is not removed. As a result, the data fluctuates around the mean and sometimes even reaches values larger than one, which is unphysical and leads to invalid data points for the bead mean squared displacement $\Delta\vec{r}^{2}(t)\right>$ and further introduces bias. This, in turn, creates difficulties in extracting the rheological information and performing the inertia corrections discussed in the following section.
To overcome this problem we propose using an essentially unbiased fit to the ICF at short times with a regularized inversion algorithm such as the well known CONTIN algorithm, which, as we show substantially improves the quality of the data at very short times. The popular open-source CONTIN algorithm by Provencher [47] inverts the following expression for the field correlation function, expressed as a spectrum over exponential decays
[…]
where $\Gamma$ denotes the decay rate. This approach was developed to model the normalized decay of a DLS field-correlation function $g_{1}(t)$ of light singly scattered from a polydisperse or multimodal suspension of colloidal particles undergoing Brownian motion. The measured intensity correlation function (ICF) $g_{2}(t)$ is related to the field correlation function via the Siegert relation: $g_{2}(t)=1+\beta g_{1}^{2}(t)$[16, 47]. Since the inversion for incomplete and noisy data is ill-posed, the algorithm employs certain regularization methods to obtain a stable fit. Numerous works have discussed improvements to the fit accuracy and stability, see Scotti et al. [48] and references therein. In the present context, our aim is not to extract or to analyse the relaxation spectrum $f(\Gamma)$, but rather to use this established method to obtain an unbiased (or at least much less biased) fit to the experimental data. This fit removes experimental noise at short times, allows extrapolation to $t=0$ to accurately determine $\beta$ and the normalized ICF can subsequently be used for further processing.
@@ II.3 One-bead microrheology with inertia corrections
The time relaxation of the intensity correlation function (ICF) is determined by the motion of the beads and the light scattering process. Precise analytical formulas have been developed to extract the bead mean square displacement (MSD) from the measured DWS intensity autocorrelation functions in transmission geometry. The measured mean square displacement, in turn, is determined by the linear viscoelastic properties of the host medium and, at short times, also by the inertia of the bead and the fluid. Particle interactions, convection or flow, and phase separation can also influence the particle motion, but these effects will not be discussed here [3, 13]. Ignoring such effects, the analysis procedure proceeds as follows: First, we extract the mean squared displacement (MSD), $\langle\Delta\vec{r}^{2}(t)\rangle$, from the ICF. Next, we remove the contributions due to bead and fluid inertia to obtain the unperturbed viscoelastic spectra of the (complex fluid) host medium. The different steps and approximations made will be discussed in the following. Additionally, we will highlight the importance of clean and low-noise data to stably perform this analysis.
@@ II.3.1 Mean squared displacement
We first convert the experimental data to the MSD using Eq. (4) taken from ref. [3], assuming that the slab thickness $L$, the sample’s transport mean free path $\ell^{\ast}$ and the wavenumber $k$ are known. Since we study non-absorbing samples, we neglect the effects of absorption, but these can be easily incorporated if required [29].
$g_{2}(t)-1=\left\{\frac{\frac{L/\ell^{\ast}+4/3}{z_{0}/\ell^{\ast}+2/3}\left(\sinh\left[\frac{z_{0}}{\ell^{\ast}}\sqrt{k^{2}\langle\Delta\vec{r}^{2}(t)\rangle}\right]+\frac{2}{3}\sqrt{k^{2}\langle\Delta\vec{r}^{2}(t)\rangle}\cosh\left[\frac{z_{0}}{\ell^{\ast}}\sqrt{k^{2}\langle\Delta\vec{r}^{2}(t)\rangle}\right]\right)}{\left(1+\frac{4}{9}k^{2}\langle\Delta\vec{r}^{2}(t)\rangle\cdot\right)\sinh\left[\frac{L}{\ell^{*}}\sqrt{k^{2}\langle\Delta\vec{r}^{2}(t)\rangle}\right]+\frac{4}{3}\sqrt{k^{2}\langle\Delta\vec{r}^{2}(t)\rangle}\cosh\left[\frac{L}{\ell^{*}}\sqrt{k^{2}\langle\Delta\vec{r}^{2}(t)\rangle}\right]}\right\}^{2}$ (4)
If internal reflections are absent or neglected, $z_{0}/\ell^{\ast}\simeq 2/3$. In our case, we consider a glass cuvette in air; thus, $z_{0}/\ell^{\ast}\simeq 2$, a value derived from the work of Durian et al. [49]. The larger $L/\ell^{\ast}$, the less important the choice of $z_{0}$ becomes[13]. We note that, for each value of $t$, we obtained a precise value for $\Delta\vec{r}^{2}(t)\right>$, and there is no fitting or ambiguity involved in this process.
@@ II.3.2 Microrheology Routine
We use the following microrheology routine, based on the work by Mason [50], to determine the apparent complex modulus (inertia-affected) $Z^{\ast}(\omega)$ from the mean squared displacement data $\langle\Delta r^{2}(t)\rangle$, determined in the previous step, as follows:
$\alpha(\omega)=\left.\frac{d\ln\langle\Delta\vec{r}^{2}(t)\rangle}{d\ln t}\right|_{t=1/\omega},$ (5)
$i\omega F_{u}\{\langle\Delta\vec{r}^{2}(t)\rangle\}\approx\langle\Delta\vec{r}^{2}\left(1/\omega\right)\rangle\,\Gamma[1+\alpha(\omega)]i^{-\alpha(\omega)},$ (6)
and finally
$Z^{\ast}(\omega)=\frac{k_{B}T}{\pi a\omega F_{u}\{\langle\Delta\vec{r}^{2}(t)\rangle\}}.$ (7)
where $a$ denotes the radius of the tracer beads. Eq. (7) is also known as the generalized Stokes Einstein equation (GSE) (in the Fourier domain) [50, 28].
@@ II.3.3 Inertia correction
At very short times, the common description of thermal Brownian motion breaks down, a phenomenon that has been well-known since Einstein’s work in 1905 [51]. The transition from ballistic to Brownian dynamics, which is strongly affected by hydrodynamics, has garnered renewed attention over the past decade [31, 52, 53]. Neglecting these contributions in the microheology analysis can result in significant errors. Properly accounting for these effects at short times, corresponding to high frequencies ($\omega$) when determining the complex modulus $G^{\ast}(\omega)$, enables the extraction of critical data that would otherwise be challenging or impossible to obtain.
The inertia correction formula used in the present study is based on the work of Domínguez-García et al. [54]. The complete theoretical treatment applied here was first developed and published by Schieber et al. [55]. Prior to that, an (iterative) approximative algorithm to correct for inertia effects was developed in Willenbacher et al. [7] and similarly employed by Mizuno et al. [56]. From the apparent complex modulus $Z^{\ast}(\omega)$, extracted from the experimental data via Eq. (7), we can calculate the host medium’s complex modulus as follows:
@@
[…]
@@ III.1 Sample
We study DWS microrheology on a wormlike micellar sample discussed extensively in previous work [58, 40, 7, 59]. To this end we prepare an aqueous solution of cetylpyridinium chloride and sodium salicylate (100 mM CPyCl–60 mM NaSal) at different temperatures. We mix the sample with polystyrene beads with a diameter of 420 nm to a final concentration of slightly below 1% in volume. We prepare the sample such that the micellar concentration relative to the aqueous solvent phase is kept constant. The sample displays nearly perfect Maxwellian behavior at low frequencies or slow relaxation times and exhibits the characteristic dynamics of polymer relaxations at short times. It is a strongly viscoelastic liquid with a terminal relaxation time that reacts to temperature over a wide range. Here, we use it as a model system to demonstrate our improved implementation of two-cell echo diffusing wave spectroscopy for different terminal relaxation times larger than or comparable to $\tau^{\mathrm{TC}}$.
@@ III.2 Experimental Setup
We carry out DWS experiments in transmission geometry, where an expanded incident laser beam illuminates the sample contained in a rectangular glass cuvette, with a width larger than its thickness $L$. The diffusely scattered light is recorded on the opposite side, in transmission, using a single-mode optical fiber. We use a classical custom-made DWS setup with a Helium-Neon gas laser operating at a 632.8 nm wavelength emission. The collimated laser beam is directed towards a 12.5mm diameter holographic diffusor (Edmund Optics, UK) with a 5 diffusing angle. Compared to ground or opal glass diffusers, holographic diffusers offer higher light transmission and a more even distribution of light. They provide control over the diffusion angle, eliminating the need for an additional collimation lens between the diffuser and the sample. In our setup the diffusor is mounted on a Voice Coil Actuated Flexure Scanner (VCFL35, Thorlabs, USA). The scanner is operated at 25 Hz using a function generator. The light diffracted from the diffusor in the forward direction forms a random speckle field that we use to illuminate the sample cuvette. The sample is contained in 12.5 mm wide, 10 mm inner width glass cuvette (Hellma, Germany) with a path length of $L=2$ mm. For the chosen distance of the ground glass from the cuvette, the speckle beam size is comparable in width to the cuvette. We place a black circular aperture with an 8 mm opening in front of the cuvette to avoid illuminating the cuvette edges.
@@
The light transmitted through the multiple scattering sample is detected in the far field on the other side using a single-mode fiber (OZ Optics, Canada). The single-mode fiber is split 50/50, and each arm is connected to an APD single-photon detector (Excelitas, USA). Using a digital correlator (LS Instruments, Switzerland), we calculate the pseudo cross-correlation function of the scattered photons, which is equal to the photon autocorrelation function but eliminates detector noise (also known as afterpulsing effects). The setup is similar to the one shown in Figure 1 of reference [42].
We operate the voice coil scanner in two modes. In slow mode, the ground glass motion leads to a slow decay of the intensity correlation function (ICF) on a timescale of $\tau^{\mathrm{TC}}$ but does not affect the short-time dynamics except for providing proper ensemble averaging of the correlation function, Fig. 1 (a). We perform an approximate normalization of $g_{2}^{\mathrm{TC}}(t)-1$ by dividing the data by the values recorded around $t=10^{-7}\,\mathrm{s}$, which typically range between 0.9 and 0.95. In fast mode, the periodic oscillation of the ground glass with a period of $\tau^{\mathrm{echo}}=0.04$ sec leads to correlation echoes [43, 42, 44], Fig. 1 b), which we detect using a linear correlator scheme configured in the digital correlator. This scheme provides a time resolution of 1 $\mu$s and spans 16 channels centered around $\tau^{\mathrm{echo}}=0.04$ sec and multiples thereof. From a Lorentzian fit of the detected correlation echoes, we determined the total area of the echo peak. This total area is a robust measure of the intensity correlation function, modulo an unknown prefactor $1/A$ [43, 42].
Figure 1: (a) Two-cell correlation function $g_{2}^{\mathrm{TC}}(t)-1$ recorded over 300 seconds for a wormlike micellar solution at T$=20^{\circ}$ with polystyrene tracer beads embedded. The overlap region is marked with an orange line. The dashed line shows the correlation function measured for a solid piece of Teflon, recorded over 3 hours and following Eq. (2) with $A=1$ and $\tau_{\text{TC}}=0.708$ s. (b) Echo two-cell DWS for the same sample recorded over 60 seconds, with data up to 12 seconds. The oscillation period is $\tau_{\mathrm{Echo}}=0.04\,\mathrm{s}$. Inset: Ratio of the two-cell correlation function to the echo correlation function. Line: Best fit to the data with $A=0.797$ and $\tau_{\mathrm{TC}}=0.697\,\mathrm{s}$ using Eq. (2). From nine datasets taken at temperatures $20^{\circ}$, $25^{\circ}$ and $30^{\circ}$ we find $\langle A\rangle=0.8468\pm 0.0774$, $\langle\tau_{\mathrm{TC}}\rangle=0.681\pm 0.0673$ or a statistical variation of about 10% for both parameters. (c) Blending and merging two-cell DWS and echo data. Blue open square: Two-cell correlation function. Dashed line: Echo correlation function. Red line: Merged correlation function obtained by multiplying the echo data by the parameter $A$ and dividing the two-cell data by the exponential function, Eq. (2) (up to $t=6\times\tau_{Echo}$) then merging the two data sets.

### v2

@@ II.2 Analysis of the short-time correlation function decay
For $t\to 0$ the ICF tends to a constant value of $g_{2}(0)-1=\beta$ where $\beta\lesssim 1$ denotes the coherence factor of the detection fiber optics and $\beta+1$ corresponds to the zero-time intercept of the ICF. In a typical DWS microrheology experiment, for short relaxation times up to some tens or hundreds of microseconds, the bead motion often leads to relatively minor decays of the measured ICF, often only by a few per cent. Extracting $\left<\Delta\vec{r}^{2}(t)\right$> from these minor decays is strongly affected by the accuracy with which one can determine $\beta$ and by the significant photon shot noise at short correlation lag times [53]. Normalization to determine $\beta$ is usually performed by extrapolating the measured $g_{2}(t)$ to $t=0$ using a linear or polynomial fitting function. Once $\beta$ is known, the normalized ICF is used to determine, e.g. $\left<\Delta\vec{r}^{2}(t)\right$>. This standard procedure suffers from two problems. First, the choice of the polynomial order used for the fit, as well as the time interval used for the fit of the data, influences the exact value of $\beta$ obtained. Here, it is important to note that, due to the inertia of the beads and the fluid, the short-time dynamics is rather complex even for simple colloidal suspensions in water, and thus no simple temporal scaling of the mean squared displacement $\left<\Delta\vec{r}^{2}(t)\right$> can be assumed [31, 20, 54]. Second, since the raw ICF is normalized by $\beta$ but is used as is for further analysis, the experimental shot noise is not removed. As a result, the data fluctuate around the mean and reach values greater than one, which is unphysical and leads to invalid data points for the bead mean squared displacement $\left<\Delta\vec{r}^{2}(t)\right$>, which introduces further bias. This, in turn, creates difficulties in extracting the rheological information and performing the inertia corrections discussed in the following section.
II.2.1 Intercept-Adjusted Exponential Basis Fit
To obtain a smooth representation of the experimental data we approximate the electric field autocorrelation function $g_{1}(t)$ using a fit with fixed, logarithmically spaced decay times:
$g_{1}(t)=\sum_{i}P_{i}\,e^{-t/\alpha_{i}},$ (3)
where the basis functions are exponentials with fixed decay times $\alpha_{i}$, and the amplitudes $P_{i}\geq 0$ are determined by fitting. We employ $N$ logarithmically spaced decay times defined as
$\alpha_{i}=10^{-7}\,\mathrm{s}\,\times e^{\frac{i-1}{u}},\qquad i=1,\dots,N,$ (4)
$\alpha_{N}$ is chosen to be one or two orders of magnitude larger than the time window of interest. Solving for $u$ gives $u=(N-1)/\ln\left(\alpha_{N}/\alpha_{0}\right).$ For viscoelastic solid samples exhibiting a stable plateau, the fit can be further improved by including a constant offset term in Eq. (3).
Here we use $N=45$ and $u=2$ $\left(\alpha_{N}=360s\right)$ to ensure robustness, but tests with smaller values yield similar results at a significantly lower computational cost as long as $N\geq 10$, see Fig. 2 (c). This exponential basis approach provides a stable and flexible reconstruction of $g_{1}(t)$, enabling an accurate representation even at very short times. Applying the Siegert relation, the intensity correlation function is modeled as:
$g_{2}(t)-1=\left(\sum_{i}P_{i}e^{-t/\alpha_{i}}\right)^{2}$ (5)
This formulation implicitly adjust the intercept via
$\beta+1=g_{2}(0)=1+\left(\sum_{i}P_{i}\right)^{2}$ (6)
and thus allows us to normalize the ICF, avoiding the need for separate normalization or intercept fitting. The fitting procedure minimizes the mean squared deviation between the modeled and measured ICF:
$\text{Error}=\frac{1}{{N^{\prime}}}\sum_{j=1}^{N^{\prime}}\left[\left(\sum_{i}P_{i}e^{-t_{j}/\alpha_{i}}\right)^{2}-\left(g_{2}(t_{j})-1\right)\right]^{2}$ (7)
where $j=1\to N^{\prime}$ is the data-point index in the multiple-tau channel layout of the ICF, and $N^{\prime}$ is the total number of data points.
After obtaining the optimal weights $P_{i}$, the model is used to reconstruct a smoothed, intercept-corrected ICF:
$g_{2}^{\text{fit}}(t)-1=\left(\sum_{i}P_{i}e^{-t/\alpha_{i}}\right)^{2}/\left(\sum_{i}P_{i}\right)^{2}$ (8)
Finally, the MSD is extracted by numerically inverting this fitted correlation function using the full Diffusing Wave Spectroscopy (DWS) model in transmission geometry, as given in Eq. (9).
We emphasize that this fitting procedure does not constitute an additional smoothing step in the DWS microrheology analysis; rather, it replaces the conventional smoothing fit applied to the mean-square displacements [25], offering the advantages discussed above.
@@ II.3 One-bead microrheology with inertia corrections
The time relaxation of the intensity correlation function (ICF) is determined by the motion of the beads and the light scattering process. Precise analytical formulas have been developed to extract the bead mean square displacement (MSD) from the measured DWS intensity autocorrelation functions in transmission geometry. The measured mean square displacement, in turn, is determined by the linear viscoelastic properties of the host medium and, at short times, also by the inertia of the bead and the fluid. Particle interactions, convection or flow, and phase separation can also influence the particle motion, but these effects will not be discussed here [3, 13]. Ignoring such effects, the analysis procedure proceeds as follows: First, we extract the mean squared displacement (MSD), $\langle\Delta\vec{r}^{2}(t)\rangle$, from the ICF. Next, we remove the contributions due to bead and fluid inertia to obtain the unperturbed viscoelastic spectra of the (complex fluid) host medium. The different steps and approximations made will be discussed below. In addition, we will highlight the importance of clean and low-noise data to stably perform this analysis.
@@ II.3.1 Mean squared displacement
We first convert the experimental data to the MSD using Eq. (9) taken from ref. [3], assuming that the slab thickness $L$, the sample’s transport mean free path $\ell^{\ast}$, and the wavenumber $k$ are known. Since we study non-absorbing samples, we neglect the effects of absorption, but these can be easily incorporated if required [29].
$g_{2}(t)-1=\left\{\frac{\frac{L/\ell^{\ast}+4/3}{z_{0}/\ell^{\ast}+2/3}\left(\sinh\left[\frac{z_{0}}{\ell^{\ast}}\sqrt{k^{2}\langle\Delta\vec{r}^{2}(t)\rangle}\right]+\frac{2}{3}\sqrt{k^{2}\langle\Delta\vec{r}^{2}(t)\rangle}\cosh\left[\frac{z_{0}}{\ell^{\ast}}\sqrt{k^{2}\langle\Delta\vec{r}^{2}(t)\rangle}\right]\right)}{\left(1+\frac{4}{9}k^{2}\langle\Delta\vec{r}^{2}(t)\rangle\right)\sinh\left[\frac{L}{\ell^{*}}\sqrt{k^{2}\langle\Delta\vec{r}^{2}(t)\rangle}\right]+\frac{4}{3}\sqrt{k^{2}\langle\Delta\vec{r}^{2}(t)\rangle}\cosh\left[\frac{L}{\ell^{*}}\sqrt{k^{2}\langle\Delta\vec{r}^{2}(t)\rangle}\right]}\right\}^{2}$ (9)
If internal reflections are absent or neglected, $z_{0}/\ell^{\ast}\simeq 2/3$. In our case, we consider a glass cuvette in air; thus, $z_{0}/\ell^{\ast}\simeq 2$, a value derived from the work of Durian et al. [55]. The larger $L/\ell^{\ast}$, the less important the choice of $z_{0}$ becomes [13]. By numerical inversion of Eq. (9) for each value of $t$, we obtain a precise value of $\left<\Delta\vec{r}^{2}(t)\right$> for each value of $g_{2}^{\text{fit}}(t)-1$, with no fitting or ambiguity involved in the process.
@@ II.3.2 Microrheology Routine
We use the following microrheology routine, based on Mason’s work [56], to determine the apparent complex modulus (inertia affected) $Z^{\ast}(\omega)$ from the mean squared displacement data $\langle\Delta r^{2}(t)\rangle$, determined in the previous step, as follows:
$\alpha(\omega)=\left.\frac{\rm{d}\ln\langle\Delta\vec{r}^{2}(t)\rangle}{\rm{d}\ln t}\right|_{t=1/\omega},$ (10)
$i\omega F_{\omega}\{\langle\Delta\vec{r}^{2}(t)\rangle\}\approx\langle\Delta\vec{r}^{2}\left(\frac{1}{\omega}\right)\rangle\,\Gamma[1+\alpha(\omega)]i^{-\alpha(\omega)},$ (11)
where $t$ is substituted by $1/\omega$, and finally
$Z^{\ast}(\omega)=\frac{k_{B}T}{\pi a\omega F_{\omega}\{\langle\Delta\vec{r}^{2}(t)\rangle\}}.$ (12)
where $a$ denotes the radius of the tracer beads. Eq. (12) is also known as the generalized Stokes-Einstein equation (GSE) (in the Fourier domain) [56, 28].
We note that Evans et al. developed a direct conversion algorithm from the mean-squared displacement (MSD)—or equivalently, from the creep compliance (see, e.g., Eq. (3.99) in [28])—to the storage and loss moduli [57, 58]. This method circumvents the approximation used in Eq. (11). It was later refined by Tassieri et al. [58]. Figure S4 shows that our approach and the direct conversion yield nearly identical results. The method by Evans et al. tends to produce noisier data, while the refined version by Tassieri et al. mitigates this issue by avoiding an approximation that can introduce artificial scatter [59]. Small residual differences remain, but a detailed analysis of these discrepancies lies beyond the scope of the present work and should be addressed in future studies.
@@ II.3.3 Inertia correction
At very short times, the common description of thermal Brownian motion breaks down, a phenomenon that has been well known since Einstein’s pioneering work on the subject in 1905 [60]. The transition from ballistic to Brownian dynamics, which is strongly affected by inertia and hydrodynamics, has attracted renewed attention more recently [61, 62, 63, 64]. Neglecting these contributions in the microheology analysis can result in significant errors. Properly accounting for these effects at short times, corresponding to high frequencies ($\omega$) when determining the complex modulus $G^{\ast}(\omega)$, enables the extraction of critical data that would otherwise be challenging or impossible to obtain.
The inertia correction formula used in the present study, Eq. (13), is based on the work of Domínguez-García et al. [65]. The complete theoretical treatment applied here was first developed and published by Schieber et al. [66]. Prior to that, an (iterative) approximative algorithm to correct for inertia effects was developed by one of us in Willenbacher et al. [7] and similarly employed by Mizuno et al. [67]. From the apparent complex modulus $Z^{\ast}(\omega)$, extracted from the experimental data, we calculate the complex modulus of the host medium as follows:
@@
where $\rho_{b}$ is the bead density and $\rho$ is the medium density in $\text{kg/m}^{3}$. The resulting $G^{\ast}(\omega)$ is a complex functions, which can be evaluated using computational tools such as Mathematica (Wolfram, USA) [68] or Python (Python Software Foundation) [69]. We can then obtain the following:
@@ III.1 Sample
We study DWS microrheology on a worm-like micellar sample discussed extensively in previous work [70, 41, 7, 71]. To do this, we prepare an aqueous solution of cetylpyridinium chloride and sodium salicylate (100 mM CPyCl - 60 mM NaSal). We mix the sample with polystyrene beads with a diameter of 420 nm to a final concentration slightly below 1% in volume. We determine the transport mean free path $\ell^{\ast}$ using the commercial ’DWS RheoLab’ instrument (LS Instruments, Switzerland) [29], and, using Mie scattering theory, correct the obtained value for the slight mismatch in laser wavelength. We prepare the sample such that the micellar concentration relative to the aqueous solvent phase is kept constant. The sample displays nearly perfect Maxwellian behavior at low frequencies or slow relaxation times and exhibits the characteristic dynamics of polymer relaxations at short times. It is a strongly viscoelastic liquid with a terminal relaxation time that reacts to temperature over a wide range. Here, we use it as a model system to demonstrate our improved implementation of two-cell echo diffusing wave spectroscopy for different terminal relaxation times larger than or comparable to $\tau^{\mathrm{TC}}$.
@@ III.2 Experimental Setup
We carry out DWS experiments in transmission geometry, where an expanded incident laser beam illuminates the sample contained in a rectangular glass cuvette, with a width larger than its thickness $L$. The diffusely scattered light is recorded on the opposite side, in transmission, using a single-mode optical fiber. We use a classical custom-made DWS setup with a Helium-Neon gas laser operating at a 632.8 nm wavelength emission. The collimated laser beam is directed towards a 12.5mm diameter holographic diffusor (Edmund Optics, UK) with a 5 diffusing angle. Compared to ground- or opal-glass diffusers, holographic diffusers offer higher light transmission and a more even distribution of light. They provide control over the diffusion angle, eliminating the need for an additional collimation lens between the diffuser and the sample. In our setup the diffusor is mounted on a Voice Coil Actuated Flexure Scanner (VCFL35, Thorlabs, USA). The scanner is operated at 25 Hz using a function generator. The light diffracted from the diffuor in the forward direction forms a random speckle field that we use to illuminate the sample cuvette. The sample is contained in a 12.5 mm wide (10 mm inner width) glass cuvette (Hellma, Germany) with a path length of $L=2$ mm. For the distance chosen for the ground glass from the cuvette (a few centimeters), the speckle beam size is comparable in width to that of the cuvette. We place a black circular aperture with an 8 mm opening in front of the cuvette to avoid illuminating the edges of the cuvette.
[…]
@@
The light transmitted through the multiple scattering sample is detected in the far field on the other side using a single-mode fiber (OZ Optics, Canada). The single-mode fiber is split 50/50 and each arm is connected to an APD single-photon detector (Excelitas, USA). Using a digital correlator (LS Instruments, Switzerland), we calculate the pseudo-cross-correlation function of the scattered photons, which is equal to the photon autocorrelation function, but eliminates detector noise (also known as afterpulsing effects). The setup is similar to the one shown in Figure 1 of reference [43].
[…]